OpenStudy (anonymous):

1.)A car travelling at 15m/s , slows down uniformly at rest in 6.0 a)how far does the car travel during the 6.0 s seconds? b)what is the car travel during the 6.0 seconds? 2.)A car takes 10 s to go from 0m/s to 24 m/s at constant acceleration a)what is the value of a b)find the distance for this

OpenStudy (anonymous):

for the 1st one i can do distance like what we used earlier right

OpenStudy (isaiah.feynman):

We need acceleration first. Its not 9.8 this time.

OpenStudy (anonymous):

so since velocity is change i velocity/ time we use this eq

OpenStudy (isaiah.feynman):

Okay we will use this equation. \[D=\frac{ 1 }{ 2 }(u+v)t \]Now can you show me how to use that equation?

OpenStudy (anonymous):

yes:)

OpenStudy (isaiah.feynman):

Okay what is the initial velocity?

OpenStudy (anonymous):

ok would it be 1/2(1+6)

OpenStudy (anonymous):

initial v is 15

OpenStudy (isaiah.feynman):

Yes, and final velocity?

OpenStudy (anonymous):

6.0

OpenStudy (isaiah.feynman):

6.0 is time (In the question). Final velocity is 0 "slows down uniformly at rest"

OpenStudy (anonymous):

kk initial velocity is 15 and the final is 0 so 1/2 (15+0)6.0

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

k i got 45

OpenStudy (isaiah.feynman):

Correct!

OpenStudy (anonymous):

yay:)

OpenStudy (anonymous):

so for be all i need to do is d/t right which is 15/6.0?

OpenStudy (isaiah.feynman):

Brb gotta do stuff in the house.

OpenStudy (anonymous):

kk

OpenStudy (isaiah.feynman):

I don't quite understand the 1b

OpenStudy (anonymous):

oh its suppose to say what is the cars average velocity during 6.0 seconds .i got 2.5

OpenStudy (isaiah.feynman):

The average velocity is the distance/time= 45/6=7.5m/s

OpenStudy (anonymous):

smt yeah i got that t dnt know why i typed that

OpenStudy (isaiah.feynman):

Its okay.

OpenStudy (anonymous):

for 2 a its acceeration so i can use vf-vi/t?

OpenStudy (isaiah.feynman):

That's what you use.

OpenStudy (anonymous):

im right??

OpenStudy (isaiah.feynman):

Yes you are.

OpenStudy (anonymous):

yay i got 25-0/10=2.5

OpenStudy (isaiah.feynman):

24 not 25 so a=2.4

OpenStudy (anonymous):

how come i get 2.5 in my calculator?

OpenStudy (isaiah.feynman):

Just a typo.

OpenStudy (anonymous):

kk so if i fint distance it would be 1/2 at^2

OpenStudy (isaiah.feynman):

Yes. Note that a=2.4

OpenStudy (anonymous):

y do i keep geting 2.5 isaiah???:(

OpenStudy (isaiah.feynman):

The question says "A car takes 10 s to go from 0m/s to 24 m/s at constant acceleration"

OpenStudy (anonymous):

1/2 (5)(0)(25)

OpenStudy (isaiah.feynman):

1/2(2.5)(100)

OpenStudy (anonymous):

y u did use 10 s instead of 100?

OpenStudy (isaiah.feynman):

t^2. t is 10, t^2 is 10^2=100

OpenStudy (anonymous):

ohhhhhhhh k i get it

OpenStudy (anonymous):

k thanks alot u r very patient

OpenStudy (isaiah.feynman):

Thank you.

OpenStudy (isaiah.feynman):

What's your name?

OpenStudy (anonymous):

rissa

OpenStudy (isaiah.feynman):

Schon, dich kennen zu lernen Rissa. :)

OpenStudy (anonymous):

what does that mean? nic e to meet you rissa?

OpenStudy (isaiah.feynman):

Yes. You know German?

OpenStudy (anonymous):

lol nope